# SlideSLAM: Sparse, Lightweight, Decentralized Metric-Semantic SLAM for Multi-Robot Navigation

### v5

@@ III-B Dec-Metric-Semantic SLAM for multi-robot exploration
In this section, we lay down the general active decentralized metric-semantic SLAM problem (Dec-Metric-Semantic SLAM). Let $\mathcal{I}_{t}^{(k)}$ be all past observations that robot $k$ has accrued by time $t$, both from its own measurements, as well as from messages it has received from other robots. At time $t$, every robot $k$ executes an exploration or navigation policy:
@@
In our metric-semantic SLAM formulation, the quantities of interest include the robot poses $\mathbf{x}_{t}^{(k)}$ and the map of the environment. At each time $t$, robot $k$ receives odometry measurements
@@
We assume $(\mathbf{w}_{t}^{(k)})_{t\leq T,k\leq K}$ and $(\mathbf{v}_{t}^{(k)})_{t\leq T,k\leq K}$ are independent random variables with zero mean and covariance $\Lambda_{k}$ and $\Gamma_{k}$, respectively.
@@
where $[K]:=\{1,\ldots,K$. For any distinct pair of robots $(i,j)$, we denote their last meeting time up to time $t$ as
@@
noting the symmetry in the superscript, i.e., $L_{t}^{(i,j)}=L_{t}^{(j,i)}$. Then, the information state of robot $k$ may be defined recursively via
$\mathcal{I}_{t}^{(k)}=\bigcup_{j\in[K]\setminus\{k\}}\mathcal{I}_{L_{t}^{(k,j)}}\ \cup\ \bigcup_{s\in[t]}\{\mathbf{h}_{t}^{(k)}\}\ \cup\ \bigcup_{s\in[t]}\{\mathbf{m}_{t}^{(k)}\},$ (7)
with the initial condition $\mathcal{I}_{0}^{(k)}=\emptyset$ for all $k\in[K]$.
We focus on a solution that uses a fixed set of policies $\mu$ that take as input compressed representations of $\mathcal{M}$ using a custom data structure, i.e., a set of object shape models with corresponding class labels, as well as a custom rule for updating that data structure, both in light of new measurements from the robot, and the robots it communicates with. Therefore, the exploration or navigation policy of robot $k$ may then be stated as
$\mu_{t}^{(k)}(\mathcal{I}_{t}^{(k)})=\mu_{t}^{(k)}(\Pi(\mathcal{I}_{t}^{(k)})),$ (8)
@@
$\Pi$ extracts an estimate of the previous trajectory of the robot, as well as its estimate of the map via
$\hat{\mathbf{x}}_{1:t}^{(k)},\hat{\mathcal{M}}_{t}^{(k)}=\underset{\mathbf{x}_{1:t}^{(k)},\mathcal{M}}{\text{argmax}}\ P(\mathbf{x}_{1:t}^{(k)},\mathcal{M}\ |\ \mathcal{I}_{t}^{(k)}).$ (10)
Finally, we use the following form of messages
@@
In principle, the robot $j$ can construct ^$\hat{\mathcal{M}}_{t}^{(i)}$ from $\mathcal{I}_{t}^{(i)}$. Given the fact that the compressed representation ^$\hat{\mathcal{M}}_{t}^{(i)}$ we use is memory efficient, we directly share this information to avoid the extra computation brought about by the reconstruction step. ^$\hat{\mathcal{M}}_{t}^{(i)}$ is shared only for checking inter-robot loop closures and estimating the transformations. Once a valid inter-robot loop closure is established, $\mathcal{I}_{t}^{(i)}$ is used to form the updated $\mathcal{I}_{t}^{(j)}$ as mentioned in Eq. 7. $\mathcal{I}_{t}^{(j)}$ will then be used to update ^$\hat{\mathcal{M}}_{t}^{(j)}$.
@@ IV-A Approach Overview
We next turn to a high-level overview of the approach for computing the information projection operator $\Pi$, which is one of the core contributions of this paper. Unlike traditional geometric-only methods such as point cloud registration, our method leverages both geometric and semantic information to find common components of maps estimated by different robots, and then refines such estimates using both its own measurements and measurements made by other robots. We illustrate this idea with a pair of robots $k$ and $j$ at time $t$, though the argument can be readily generalized to interactions with a larger number of robots simultaneously. It solves the maximum likelihood state estimation problem as follows. For every robot $j\in\mathcal{N}_{t}^{(k)}$, we consider the message $r_{j\rightarrow k}(t)=(\mathcal{I}_{t}^{(j)},\hat{\mathcal{M}}_{t}^{(j)})$. First we determine common map observations made by robots $k$ and $j$ up to time $t$ by performing inter-robot loop closure detection on ^$\hat{\mathcal{M}}_{t}^{(j)}$ and ^$\hat{\mathcal{M}}_{t}^{(k)}$. The output of the loop closure module yields an estimate of the relative transformation ${}^{(k)}T_{(j)}=({}^{(k)}R_{(j)},{}^{(k)}\mathbf{t}_{(j)})\in SE(3)$ between reference frames of robot $j$ and robot $k$. In particular a point with coordinates $p$ in the reference frame of robot $j$ has coordinates ${}^{(k)}\textbf{p}={}^{(k)}R_{(j)}{}^{(j)}\textbf{p}+{}^{(k)}\mathbf{t}_{(j)}$ in the reference frame of robot $k$. This transformation is then used to determine which landmarks in the map of robot $j$ correspond to landmarks that also exist in the map of robot $k$. Thereafter, the trajectories of both robot $j$ and $k$ are optimized together with the union of object models in the map, taking care to associate measurements of the same landmark taken by two robots with the same variable in the factor graph.
@@
Our system is illustrated in Fig. 3. Our framework works on robots with different types of sensors, including RGBD cameras and LIDARs, as detailed in Section V-A. Our metric-semantic SLAM framework handles such heterogeneity by using different front-end pipelines. We integrate the proposed metric-semantic SLAM framework with our autonomous exploration and navigation stack as detailed in Section V-B. In the rest of this section, we will provide details on individual modules of the proposed framework.
@@ IV-B Map representation
Therefore, for the high-level map, we design a generic, sparse, and informative metric-semantic map representation. It contains a set of objects, each represented by a semantic class and a state vector that describes its model parameters, including position, orientation, and shape information. Such high-level maps, as illustrated in Fig. 2, can be maintained throughout the entire mission of the robot. An optional mid-level map consisting of semantically segmented point clouds will be used only upon the user’s request to avoid unnecessary memory burden. The low-level volumetric map, as displayed in Fig. 5, is designed to be computationally and memory efficient, as detailed in [6].
The high-level abstraction enables tasks including localization, mapping, map sharing, place recognition, map merging, and active exploration over a large scale and in real time. The mid-level abstraction is not required for real-time operations. However, when requested, it can be used to generate more detailed models to facilitate downstream tasks such as manipulation or detailed measurements of objects. The low-level abstraction is used for autonomous navigation and geometric exploration.
The metric-semantic SLAM problem can be broken down into two subproblems: (a) determining the discrete semantic labels and data association of detected objects and (b) optimizing over the continuous variables of robot poses and object model parameters (pose and dimension). In this work, we approach problem (a) using deep neural networks for detection and assignment algorithms for data association, and problem (b) by first converting object-level observations into factors in a factor graph and then using the incremental smoothing and mapping algorithm (iSAM2) [54, 55] to optimize it. In the rest of this section, we will provide details on these two steps.
@@ IV-C Object detection and modeling
The front-end of our SLAM framework is responsible for processing raw data from different sensors and converting it into object-level observations. The process is broken down into three components. 1) Object detection or instance segmentation. 2) Object instance tracking to accumulate observations from different views. 3) Shape model fitting to the instances based on the class labels.
IV-C1 Closed-set object detection
@@
For instance segmentation on LiDAR, we use the RangeNet++ [56] deep semantic segmentation model. This model performs segmentation on the range image, which is a spherical projection of a point cloud. This operation drastically decreases the inference time compared to performing inference on a 3D data structure such as a point cloud and makes real-time inference onboard SWaP-constrained platforms possible. We use a custom open-source labeling tool to generate training data [57]. For the RGBD sensor, we use YOLOv8 [58] to perform instance segmentation on RGB images and backproject the pixels in each object instance’s segmented mask into point clouds using depth information. We apply depth-based thresholding within each object instance’s mask, keeping only the points within a certain range of depth percentile, to avoid including noisy pixels that may be falsely included in the mask.
IV-C2 Open-vocabulary object detection
The main limitation of the above-mentioned object detection or instance segmentation neural networks is that they only enable the robot to perceive a subset of the semantic information in the world characterized by pre-identified labels. Our goal is to go beyond such closed-set object detection and semantic segmentation, and enable the robot to perceive open-vocabulary semantic information, such that the robot can construct a metric-semantic map that allows it to interact and accomplish tasks in open-world settings. We achieve this by incorporating open-vocabulary object detectors into our system. Open-vocabulary object detectors, such as [59, 60, 61], are trained on large-scale (often internet-scale) image-text datasets. This allows these models to achieve zero-shot transfer to novel environments with a wide range of object categories. In addition, such models take in natural language prompts, instead of being limited to a set of pre-identified labels. Such capabilities enable robots to operate in novel environments, accomplish novel tasks, and even come up with task-relevant prompts to guide their navigation and exploration.
@@ IV-C2 Open-vocabulary object detection
We utilize the YOLO-World model [61] as our open-vocabulary object detector due to its real-time onboard performance. The YOLO-World model processes pairs of images and a list of text prompts, where we specify a set of object categories of interest in the text prompts. The model then outputs object detections corresponding to each query in the text prompts. We apply depth thresholding to these detections, similar to the approach used for YOLOv8. Our system can also be easily adapted to integrate other open-vocabulary object detectors that output bounding boxes of objects.
IV-C3 Object modeling
@@
Once the object instances are extracted, these objects are tracked over time using the Hungarian assignment algorithm [62]. Tracking instances across time and from different viewpoints ensures robust geometry of objects and minimizes false positives by rejecting those tracks that only appeared once or very few times, which is necessary for metric-semantic mapping. Finally, the algorithm also helps with robustness against moving objects by tracking and accumulating their point clouds. For example, an accumulated instance of a moving object appears elongated when compared to its true size, and such objects can then be filtered out by simply applying a threshold on their dimensions. This preliminary tracking step is used to more robustly obtain object-level observations. In the factor graph optimization back end, there is a separate object tracking step that associates these object observations with existing object landmarks to construct factors in the factor graph.
@@ IV-D1 Cuboid factors
The treatment of cuboid factors is the same as presented in our previous work [3]. We first recover the cuboid’s $\mathbb{SE}$(3) pose based on its state vector $\ell_{i}^{g}=[\mathbf{r};\mathbf{t};\mathbf{d}]$, and transform it from the reference frame to the body frame using ${\mathbf{H}^{s}_{\texttt{cub}}}=\mathbf{H}^{s}_{w}\mathbf{H}^{w}_{\texttt{cub}}$. The error function for cuboidal objects is as follows:
$\mathbf{e}_{\texttt{cub}}=\begin{bmatrix}\mathbf{log}((\mathbf{H}^{s}_{\texttt{cub}}(\mathbf{z}))^{-1}({\mathbf{H}^{s}_{w}}\mathbf{H}^{w}_{\texttt{cub}}))^{\vee}\\ \mathbf{d}-\mathbf{d}{(\mathbf{z})}\end{bmatrix}$ (13)
@@ IV-D2 Cylinder factors
$\mathbf{e}_{\texttt{cyl}}=\begin{bmatrix}(\mathbf{R}^{s}_{w}\mathbf{b}+\mathbf{t}^{s}_{w})-\mathbf{b}{(\mathbf{z})}\\ \mathbf{R}^{s}_{w}\mathbf{n}-\mathbf{n}{(\mathbf{z})}\\ {r}-{r}{(\mathbf{z})}\end{bmatrix}$ (14)
@@ IV-D3 Ellipsoid factors
$\mathbf{e}_{\texttt{ellip}}=\begin{bmatrix}rg_{\text{exp}}-rg{(\mathbf{z})}\\ \theta_{\text{exp}}-\theta{(\mathbf{z})}\\ \phi_{\text{exp}}-\phi{(\mathbf{z})}\\ \end{bmatrix}$ (15)
@@ IV-E Place recognition and loop closure
The two example algorithms exhibit different behaviors and offer complementary performance across various scenarios. Generally, we recommend using SlideGraph as the primary algorithm, with SlideMatch serving as an alternative option, as further discussed in detail in Section V-L.

### v6

@@ III-B Dec-Metric-Semantic SLAM for multi-robot exploration
In this section, we state the general active decentralized metric-semantic SLAM problem (Dec-Metric-Semantic SLAM). Let $\mathcal{I}_{t}^{(k)}$ be all past observations that robot $k$ has accrued by time $t$, both from its own measurements, as well as from messages it has received from other robots. At time $t$, every robot $k$ executes an exploration or navigation policy:
@@
In our metric-semantic SLAM formulation, the quantities of interest include the robot poses $\mathbf{x}_{t}^{(k)}$ and the map of the environment $\mathcal{M}$. At each time $t$, robot $k$ receives odometry measurements
@@
We assume $(\mathbf{w}_{t}^{(k)})_{t\leq T,k\leq K}$ and $(\mathbf{v}_{t}^{(k)})_{t\leq T,k\leq K}$ are independent random variables with zero mean and covariance $\Lambda_{k}$ and $\Gamma_{k}$, respectively. Currently, both variables $\mathbf{w}_{t}^{(k)}$ and $\mathbf{v}_{t}^{(k)}$ are set to constant values since key poses $\mathbf{x}_{t}^{(k)}$ are sampled at a fixed robot travel distance interval. However, in our implementation, we provide interface to allow scaling $\mathbf{w}_{t}^{(k)}$ according to travel distance and modify $\mathbf{v}_{t}^{(k)}$ accordingly to account for sensor and object detection noise.
@@
where $[K]:=\{1,\ldots,K\}$. For any distinct pair of robots $(i,j)$, we denote their last meeting time up to time $t$ as
@@
noting the symmetry in the superscript, i.e., $L_{t}^{(i,j)}=L_{t}^{(j,i)}$. Then, the information state of robot $k$ may be defined recursively as
$\mathcal{I}_{t}^{(k)}=\bigcup_{j\in[K]\setminus\{k\}}\mathcal{I}_{L_{t}^{(k,j)}}\ \cup\ \bigcup_{s\in[t]}\left(\{\mathbf{h}_{s}^{(k)}\}\ \cup\ \{\mathbf{m}_{s}^{(k)}\}\right),$ (7)
with initial condition $\mathcal{I}_{0}^{(k)}=\emptyset$ for all $k\in[K]$.
We develop an approach that uses a set of policies $\mu$ (e.g., active SLAM, exploration, etc.) that take as input compressed representations of $\mathcal{M}$ using a custom data structure, i.e., a set of object landmarks with corresponding poses, shapes, and class labels, as well as a custom rule for updating the data structure both using new measurements from the robot and information from its communication with other robots. The exploration or navigation policy of robot $k$ may then be defined via
$\mu_{t}^{(k)}(\mathcal{I}_{t}^{(k)}):=\mu_{t}^{(k)}(\Pi(\mathcal{I}_{t}^{(k)})),$ (8)
@@
$\Pi$ extracts an estimate of the previous trajectory of robot $k$, as well as its estimate of the map via
$\hat{\mathbf{x}}_{1:t}^{(k)},\hat{\mathcal{M}}_{t}^{(k)}=\underset{\mathbf{x}_{1:t}^{(k)},\mathcal{M}}{\text{argmax}}\ P(\mathbf{x}_{1:t}^{(k)},\mathcal{M}\ |\ \mathcal{I}_{t}^{(k)}),$ (10)
where $P(\mathbf{x}_{1:t}^{(k)},\mathcal{M}\ |\ \mathcal{I}_{t}^{(k)})$ represents the joint probability distribution of the robot $k$’s trajectory from the start up to time $t$, and the map $\mathcal{M}$, which includes object landmarks, conditioned on the information state $\mathcal{I}_{t}^{(k)}$ of the robot $k$. Finally, we use the following form of messages
@@
In principle, robot $j$ can construct ^$\hat{\mathcal{M}}_{t}^{(i)}$ from $\mathcal{I}_{t}^{(i)}$. Given the fact that the compressed representation ^$\hat{\mathcal{M}}_{t}^{(i)}$ we use is memory efficient, we directly share this information to avoid the extra computation brought about by the reconstruction step. The map ^$\hat{\mathcal{M}}_{t}^{(i)}$ is shared only for checking inter-robot loop closures and estimating the transformations. Once a valid inter-robot loop closure is established, $\mathcal{I}_{t}^{(i)}$ is used to form the updated $\mathcal{I}_{t}^{(j)}$ as mentioned in Eq. 7. $\mathcal{I}_{t}^{(j)}$ will then be used to update ^$\hat{\mathcal{M}}_{t}^{(j)}$.
@@ IV-A Approach overview
We next turn to an overview of the approach for computing the information projection operator $\Pi$, which is one of the core contributions of this paper. Unlike traditional geometric-only methods such as point cloud registration, our method leverages both geometric and semantic information to find common components of maps estimated by different robots, and then refines such estimates using both its own measurements and measurements made by other robots. We illustrate this idea with a pair of robots $k$ and $j$ at time $t$, though the argument can be readily generalized to simultaneous interactions of a larger number of robots. Robot $k$ solves the maximum likelihood state estimation problem at time $t$ as follows. For every robot $j\in\mathcal{N}_{t}^{(k)}$, we consider the message $r_{j\rightarrow k}(t)=(\mathcal{I}_{t}^{(j)},\hat{\mathcal{M}}_{t}^{(j)})$. First we determine common map observations made by robots $k$ and $j$ up to time $t$ by performing inter-robot loop-closure detection on ^$\hat{\mathcal{M}}_{t}^{(j)}$ and ^$\hat{\mathcal{M}}_{t}^{(k)}$. The output of the loop-closure module yields an estimate of the relative transformation ${}^{(k)}T_{(j)}=({}^{(k)}R_{(j)},{}^{(k)}\mathbf{t}_{(j)})\in\mathbb{SE}(3)$ between reference frames of robot $j$ and robot $k$. In particular, a point with coordinates ${}^{(j)}\textbf{p}$ in the reference frame of robot $j$ has coordinates ${}^{(k)}\textbf{p}={}^{(k)}R_{(j)}{}^{(j)}\textbf{p}+{}^{(k)}\mathbf{t}_{(j)}$ in the reference frame of robot $k$. This transformation is then used to determine which landmarks in the map of robot $j$ correspond to landmarks that also exist in the map of robot $k$. Thereafter, the trajectories of both robot $j$ and $k$ are optimized together with the union of object models in the map, taking care to associate measurements of the same landmark taken by two robots with the same variable in the factor graph:
@@
Our system works on robots with different types of sensors, including RGBD cameras and LIDARs, as illustrated in Fig. 3 and detailed in Section V-A. Our metric-semantic SLAM framework handles such heterogeneity by using different front-end pipelines. We integrate the proposed metric-semantic SLAM framework with our autonomous exploration and navigation stack as detailed in Section V-B. In the rest of this section, we will provide details on individual modules of the proposed framework.
@@ IV-B Map representation
Therefore, we design a generic, sparse, and informative metric-semantic map representation. It contains a set of objects, each represented by a semantic class and a state vector that describes its model parameters, including position, orientation, and shape information. Such a map representation, as illustrated in Fig. 2, can be maintained throughout the entire mission of the robot while enabling tasks such as localization, mapping, map sharing, place recognition, map merging, and active exploration over a large scale and in real time.
Figure 5: Semantics-in-the-loop autonomy. The Falcon 250 UAV autonomously explores the environment while constructing a voxel map utilized for low-level navigation and a metric-semantic map used for active semantic loop closure in real time. The long-horizon exploration trajectory (thin green line) is shown alongside the dynamically feasible trajectory (thick blue curve) planned toward the next exploration waypoint.
The metric-semantic SLAM problem can be broken down into two subproblems: (a) determining the discrete semantic labels and data association of detected objects and (b) optimizing over the continuous variables of robot poses and object model parameters (pose and shape). In this work, we approach problem (a) using deep neural networks for detection and assignment algorithms for data association, and problem (b) by first converting object observations into factors in a factor graph and then using an incremental smoothing and mapping algorithm (iSAM2) [60, 61] to optimize it. In the rest of this section, we provide details on these two steps.
@@ IV-C Object detection and modeling
The front-end of our SLAM framework is responsible for processing raw data from different sensors and converting it into object-level observations. The process is broken down into three components: 1) object detection or instance segmentation; 2) object instance tracking to accumulate observations from different views; 3) shape model fitting to the instances based on the class labels.
@@
For semantic segmentation on point cloud from LiDAR, we trained RangeNet++ [62] with a small modified backbone. This model performs segmentation on the range image, which is a spherical projection of a point cloud. This operation drastically decreases the inference time compared to performing inference on a 3D data structure such as a point cloud and makes real-time inference onboard SWaP-constrained platforms possible. We use a custom open-source labeling tool to generate training data [63].
For the RGBD sensor, we use YOLOv8 [64] to perform instance segmentation on RGB images and backproject the pixels in each object instance’s segmented mask into point clouds using depth information. We apply depth-based thresholding within each object instance’s mask, keeping only the points within a certain range of depth percentile, to avoid including noisy pixels.
@@
In addition to supporting the aforementioned closed-set object detection methods that only enable the robot to perceive a subset of the semantic information in the world characterized by pre-identified labels, we provide support to incorporate open-vocabulary object detectors into our system. Specifically, we utilize the YOLO-World model [65] as our open-vocabulary object detector due to its real-time onboard performance. The YOLO-World model processes pairs of images and a list of text prompts, where we specify a set of object categories of interest. The model then outputs object detections corresponding to each query in the text prompts. We apply depth thresholding to these detections, similar to the approach used for YOLOv8.
@@
Once the point cloud per detection is extracted, these detected objects are tracked over time using the Hungarian assignment algorithm [66]. Tracking instances across time and from different viewpoints ensures robust geometry of objects and minimizes false positives by rejecting those tracks that only appeared once or very few times, which is necessary for metric-semantic mapping. Finally, the algorithm also helps with robustness against moving objects by tracking and accumulating their point clouds. For example, an accumulated instance of a moving object appears elongated when compared to its true size, and such objects can then be filtered out by simply applying a threshold on their dimensions. This preliminary tracking step is used to more robustly obtain object-level observations. In the factor graph optimization back end, there is a separate object tracking step that associates these object observations with existing object landmarks to construct factors in the factor graph.
@@
We highlight that we do not rely on a specific object detector. Our implementation has a modular front-end design, enabling the seamless integration of alternative object detectors in the future.
@@ IV-D1 Cuboid factors
We first recover the cuboid’s $\mathbb{SE}$(3) pose based on its state vector $\ell_{i}^{g}=[\mathbf{r};\mathbf{t};\mathbf{d}]$, and transform it from the reference frame to the body frame using ${\mathbf{H}^{s}_{\texttt{cub}}}=\mathbf{H}^{s}_{w}\mathbf{H}^{w}_{\texttt{cub}}$. The error function for cuboidal objects is as follows:
$\mathbf{e}_{\texttt{cub}}=\begin{bmatrix}\mathbf{log}((\mathbf{H}^{s}_{\texttt{cub}}(\mathbf{z}))^{-1}({\mathbf{H}^{s}_{w}}\mathbf{H}^{w}_{\texttt{cub}}))^{\vee}\\ \mathbf{d}-\mathbf{d}{(\mathbf{z})}\end{bmatrix},$ (13)
@@ IV-D2 Cylinder factors
$\mathbf{e}_{\texttt{cyl}}=\begin{bmatrix}(\mathbf{R}^{s}_{w}\mathbf{b}+\mathbf{t}^{s}_{w})-\mathbf{b}{(\mathbf{z})}\\ \mathbf{R}^{s}_{w}\mathbf{n}-\mathbf{n}{(\mathbf{z})}\\ {r}-{r}{(\mathbf{z})}\end{bmatrix},$ (14)
@@ IV-D3 Ellipsoid factors
$\mathbf{e}_{\texttt{ellip}}=\begin{bmatrix}rg_{\text{exp}}-rg{(\mathbf{z})}\\ \theta_{\text{exp}}-\theta{(\mathbf{z})}\\ \phi_{\text{exp}}-\phi{(\mathbf{z})}\end{bmatrix}.$ (15)
@@ IV-E Place recognition and loop closure
The two example algorithms exhibit different behaviors and offer complementary performance across various scenarios. Generally, we recommend using SlideGraph as the primary algorithm, with SlideMatch serving as an alternative option, as further discussed in detail in Section V-J.